Cullumber Co. has a capital structure, based on current market values, that consists of 40 percent debt, 10 percent preferred stock, and 50 percent common stock. If the returns required by investors are 11 percent, 11 percent, and 15 percent for the debt, preferred stock, and common stock, respectively, what is Cullumber’s after-tax WACC? Assume that the firm’s marginal tax rate is 40 percent. (Round final answer to 2 decimal places, e.g. 15.25%.)
Answer #1

After-tax cost of debt=11*(1-tax rate)

=11*(1-0.4)=6.6%

WACC=Respective costs*Respective weight

=(0.4*6.6)+(0.1*11)+(0.5*15)

=11.24%